Circular No. 8765

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 U3 (NEAT) R. Stoss, Astronomisches Rechen-Institut, Heidelberg, and Observato-

rio Astronomico de Mallorca, reports the recovery by J. L. Ortiz and A. Mora of P/2001 K1 (NEAT) on CCD images obtained with the 2.5-m Isaac Newton Telescope at La Palma, the object appearing asteroidal on 300-s R-band frames. The astrometric observations provided below were measured by Stoss. The indicated correction to the prediction on MPC 54169 is ΔT = −0.4 day. The recovery positions appear to confirm the proposed identification of this comet with an object found by Stoss on Palomar Sky Survey plates from 1955; the 1955 astrometry falls within 15′′ of the position predicted from an orbit linking the 2001 and 2006 observations.

Date  UT  α_{2000}  δ_{2000}  Mag.
1955 May 14.22084  14^{h}18^{m}41.^{s}81  + 6^{o}22′39′′.7
14.22778  14 18 41.50  + 6 22 38.0
14.26250  14 18 40.12  + 6 22 27.3
2006 Oct. 23.13259  4 53 18.59  +14 46 16.3  22.4
23.13672  4 53 18.52  +14 46 16.2  22.2
24.03318  4 52 59.72  +14 46 22.4  22.5
24.03727  4 52 59.64  +14 46 22.5  22.2

Orbital elements by B. G. Marsden, linking 1955 to 2006:

Epoch = 1954 Dec. 4.0 TT

T = 1954 Dec. 22.7570 TT  ω = 93.^{o}0505
e = 0.345814  Ω = 86.3172 }2000.0
q = 2.564300 AU  i = 16.6409

a = 3.919835 AU n^{o} = 0.1269996 P = 7.761 years Epoch = 2000 Oct. 23.0 TT

T = 2000 Nov. 6.8998 TT  ω = 94.^{o}6503
e = 0.357232  Ω = 84.8386 }2000.0
q = 2.470580 AU  i = 16.9125

a = 3.843656 AU n^{o} = 0.1307938 P = 7.536 years Epoch = 2008 May 14.0 TT

T = 2008 May 26.7064 TT  ω = 94.^{o}9130
e = 0.357635  Ω = 84.7534 }2000.0
q = 2.468702 AU  i = 16.9133

a = 3.843143 AU n^{o} = 0.1308200 P = 7.534 years

2006 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green